# Waterfall: Framework for Robust and Scalable Text Watermarking

### v1

@@ Abstract
Protecting intellectual property (IP) of text such as articles and code is increasingly important, especially as sophisticated attacks become possible, such as paraphrasing by large language models (LLMs) or even unauthorized training of LLMs on copyrighted text to infringe such IP. However, existing text watermarking methods are not robust enough against such attacks nor scalable to millions of users for practical implementation. In this paper, we propose Waterfall, the first training-free framework for robust and scalable text watermarking applicable across multiple text types (e.g., articles, code) and languages supportable by LLMs, for general text and LLM data provenance. Waterfall comprises several key innovations, such as being the first to use LLM as paraphrasers for watermarking along with a novel combination of techniques that are surprisingly effective in achieving robust verifiability and scalability. We empirically demonstrate that Waterfall achieves significantly better scalability, robust verifiability, and computational efficiency compared to SOTA article-text watermarking methods, and also showed how it could be directly applied to the watermarking of code.
Waterfall: Framework for Robust and Scalable Text Watermarking
@@ 1 Introduction
Adding to the challenge is the growing prevalence of generative large language models (LLMs) that may be trained on copyrighted text without permission. To enforce IP rights, content creators would need to be able to do data provenance for LLMs, i.e., prove whether their set of work had been used to train 3rd party black-box LLMs. While there have been recent works tackling this problem (Abdelnabi and Fritz, 2021; Zhang et al., 2023), they largely require intervening in the training process of the LLMs. This is unrealistic in practice, as not all LLM service providers may be cooperative due to incentive misalignment, and adversaries may also use open-source LLMs.
@@
In this paper, we propose Waterfall, the first training-free framework for robust and scalable text watermarking applicable across multiple text types (e.g., articles, code) and languages supportable by LLMs, for general text and LLM data provenance. Rather than viewing LLMs as just sources of IP infringement, we introduce the novel perspective of using LLMs’ capabilities to protect existing IP. Though simple, our training-free framework comprises several key innovations such as being the first to use LLM as paraphrasers for watermarking along with a novel combination of techniques that are surprisingly effective in achieving robust verifiability, scalability, and data provenance for LLMs, surpassing state-of-the-art text watermarking methods. In summary, our contributions are as follows:
1. We introduced a formulation of the robust and scalable text watermarking problem setting and lay out a set of desiderata to be satisfied (Section 2).
2. To tackle the challenges arising from these desiderata, we proposed Waterfall comprising novel innovations, including: (a) effective use of LLM paraphrasers to watermark existing text with IP to be protected (Section 3.1); (b) combination of vocab permutation and a new orthogonal watermarking perturbation method in token space, to achieve high scalability and robust verifiability while preserving fidelity (Section 3.3).
3. We conducted comprehensive empirical evaluations, demonstrating that Waterfall achieves significantly better scalability, robust verifiability, and computational efficiency compared to SOTA article-text watermarking methods (Section 4.1), while meeting the desiderata for a variety of applications, including for LLM data provenance of articles (Section 4.3). We also showed how Waterfall could be directly applied to the watermarking of programming code (Section 4.2).
@@ 2 Problem formulation and Desiderata
Watermarking: Client $i$ uses a watermarking operator $\mathcal{W}(\mu_{i},T_{\text{o}})\rightarrow T_{\text{w}}^{(i)}$ to produce a text $T_{\text{w}}^{(i)}$ that contains watermark $\mu_{i}$, preserves $c$, and can then be used/distributed freely.
Attacks: There are adversaries who aim to claim the IP in $T_{\text{w}}^{(i)}$ through attacks $\mathcal{A}(T_{\text{w}}^{(i)})\rightarrow T_{\text{sus}}^{(i)}$ that generate their own text $T_{\text{sus}}^{(i)}$ without the watermark $\mu_{i}$ while preserving semantic content $c$. Adversaries do not know $\mu_{i}$ but are able to perform several classes of attacks:
@@
Verification: Client $i$ can use a verification operator $\mathcal{V}(\mu_{i},T_{\text{sus}})$ to generate a score $q$ indicating the likelihood that $T_{\text{sus}}$ contains $\mu_{i}$. They can then use a setting-specific threshold $\bar{q}$ to classify $T_{\text{sus}}$ as watermarked with $\mu_{i}$ if $q\geq\bar{q}$. The operator $\mathcal{V}$ should be quick and not assume access to $T_{\text{o}}$, as in practice client $i$ may have a large set of $T_{\text{w}}$ and would need to automate the application of $\mathcal{V}$ to scan through a large set of $\{T_{\text{sus}}\}$ to identify any plagiarism (further discussion in Appendix K).
@@
1. Fidelity. The watermarked text $T_{\text{w}}$ should be semantically similar to $T_{\text{o}}$, i.e., $\mathcal{S}(T_{\text{o}},T_{\text{w}})\geq s$, where $\mathcal{S}:\mathbb{T}\crossproduct\mathbb{T}\rightarrow[0,1]$ is a user-defined fidelity metric depending on the purpose and type of text (e.g., semantic similarity score for articles, or unit tests for code) and $s$ is a setting-specific threshold. We define $\mathbb{T}_{c,s}^{\mathcal{W}}=\{T\in\mathbb{T}:\mathcal{S}(T_{\text{o}},T)\geq s\}$ as the support set of all $T_{\text{w}}$ that a watermarking operator $\mathcal{W}$ can possibly generate for $T_{\text{o}}$ with content $c$ under a $s$-fidelity setting.
2. Verifiability. The verification operator $\mathcal{V}(\mu_{i},T_{\text{w}}^{(i)})$ should have high efficacy, accounting for Type I and II errors over various thresholds $\bar{q}$. We evaluate this with AUROC computed over a test set.
@@
3. Robust verifiability. The verification operator on watermarked text after attacks $\mathcal{A}\in\mathbb{A}$, i.e., $\mathcal{V}(\mu_{i},\mathcal{A}(T_{\text{w}}^{(i)}))$, retains high verifiability. This means that the watermark should remain verifiable even after attacks, which constrains framework design. For example, the verification operator should not extract $\mu$ in any subroutine, as an attacker may use it to get $\mu$ and devise an $\mathbb{A}_{3}$ attack to overwrite it (see Section 4.1).
4. Scalability. The framework should support a large $|\mathbb{M}|$ (set of IDs) while meeting all other desiderata.
@@ 3.3 Increasing scalability with vocab permutation and orthogonal perturbation
Orthogonal perturbation: Our proposed perturbation operator $\mathcal{F}$ involves two sub-operations acting on $V_{w}$. It first maps each key $k_{p}\in\mathbb{K}_{p}$ to a unique function in a pre-defined family of orthogonal functions, and then adds the chosen perturbation function to the logits $L_{j}$ of the LLM output in $V_{w}$ space:
@@ 3.4 Waterfall Framework
Practical considerations. Waterfall is highly adaptable, i.e., it can be implemented with different LLM as paraphrasers, allowing our framework to achieve better watermarking performance and support more text types as the LLM landscape evolves. Methods like prompt engineering (Wei et al., 2022; Lin et al., 2023) and Reflexion (Shinn et al., 2023; Madaan et al., 2023) may also help to boost performance in some settings, as we demonstrate in our code watermarking experiments (Section G.2). We elaborate further on possible large-scale deployment methods of Waterfall and other practical considerations in Appendix L.
@@ 4 Experiments
Figure 4: Higher watermarking strength $\kappa$ improves verifiability and extraction accuracy. (a) Increasing $\kappa$ trades off fidelity for higher verifiability. (b) Longer token length $N$ improves verifiability. (c) Combining more pieces of text improves extraction accuracy towards $100\%$. Extraction accuracy is significantly higher than random guess accuracy of $0.003125\%$.
Figure 5: More queries improve LLM provenance verifiability. Increasing number of clients $M$ in the training dataset only slightly decreases verifiability.
@@ 4.1 Data ownership
Fidelity-verifiability. We consider the fidelity and verifiability of the schemes before adversarial attacks. Verifiability is computed as the AUROC based on varying their respective classification thresholds, i.e., the verification score threshold $\bar{q}$ for Waterfall, and bit-error rate threshold for M-bit and P-nlw.
Waterfall allows for adjustable watermarking strength to calibrate the fidelity-verifiability trade-off based on the clients’ use cases. Figure 4a shows the Pareto frontier of the trade-off. Stronger watermark strength $\kappa$ improves verifiability but also introduces larger distortions to the LLM paraphrasing process, decreasing the fidelity of watermarked text. For our experiments, we mainly used $\kappa=6$, achieving a mean AUROC of 0.992 and STS of 0.887. Even for shorter texts of just 100 tokens ($about$ 65 words), Waterfall achieves high verifiability with an AUROC of 0.98 (Figure 4b). Additional results are in Section E.4.
Note that M-bit and P-nlw were designed with only one setting, allowing for only a single fidelity-verifiability score, with mean STS scores of 0.998 and 0.942 respectively, and corresponding AUROC scores of 0.987 and 0.882. While the STS scores are high, it is expected given that the schemes only make minor edits to $T_{\text{o}}$ which would be more fragile to attacks, as we will see later. Additionally, the word replacements by M-bit and P-nlw introduced noticeable linguistic errors that are difficult to evaluate and not captured by the STS score (shown in Section E.5).
Robust verifiability. We consider the various classes of attacks $\mathbb{A}$ mentioned in Section 2. Details of the setup for each attack and additional results are in Appendix F.
@@
Scalability. As mentioned in Section 3.3, Waterfall has a large maximum scalability of $M=|\mathcal{F}_{1}|\cdot|\mathbb{V}|!\sim 10^{130274}$ based on our implementation using the Fourier perturbation function, and Llama-2 model as LLM paraphraser. In comparison, M-bit and P-nlw, have scalability dependent on the number of possible synonym replacements in any given text, which is limited by text length and varies for different text. On the c4 dataset with a mean article length of 355 words, M-bit and P-nlw can only embed a mean of 9.5 bits ($M\sim 10^{3}$) and 23.2 bits ($M\sim 10^{10}$) respectively.
@@
Metadata extraction. We also demonstrate how Waterfall could be used to embed metadata while watermarking. We consider metadata $k_{p}\in\{1,2,...,31999\}$, and the task is to extract the embedded $k_{p}$ if the text has been verified as watermarked with $\mu$. We do so by using $k_{p}$ as the frequency of the Fourier perturbation function $\mathcal{F}_{1}$, and perform extraction with the DFT. Figure 4c shows the extraction accuracy of Waterfall for different perturbation magnitudes $\kappa$. By taking multiple samples of $T_{\text{w}}$, multiple articles could be combined together to improve extraction accuracy. For $\kappa=6$, accuracy increases from 48% to 99% with only 5 pieces of text. Details are in Section E.8.
Computational costs. We note that Waterfall also has lower computational cost compared to benchmarks (Table 2). Waterfall verification can be run in parallel on a CPU, requiring only 0.035s when ran on a single 16-core CPU, which is 75x and 4237x faster than M-bit and P-nlw respectively, both which require inference using deep learning models. This is important in the context of protection of IP, e.g., where data providers may have to scan through large amount of online data for any IP infringement. Further discussion on the deployment costs are in Appendix L.
@@ 4.3 LLM data provenance of articles
Verifiability. To evaluate verifiability, we queried the fine-tuned model with the first 50 tokens of a randomly chosen abstract, and applied the verifiability operator on the next 100 generated new tokens to test for the associated watermark (details in Section H.2). Our results, presented in Figure 5, shows that Waterfall has high verifiability, reaching AUROC of 1.0 with just 100 queries to the fine-tuned LLM.
Scalability. To explore the scalability of Waterfall for data provenance, we combined the datasets of different number of clients, $M\in\{1,5,10,20,100\}$, each watermarked with their own unique ID $\mu$, and use the combined dataset for fine-tuning the adversarial model. As expected, Figure 5 shows that dealing with an aggregated dataset mixed with a larger $M$ number of different watermarks would result in a decrease in verifiability. However, our results indicate that this decrease leveled off from $M=20$ to $M=100$ and still allow for an AUROC (verifiability) of 1.0 with around 100 queries even for $M=100$, demonstrating the scalability of Waterfall to a sizable number of clients.

### v2

# EMNLP 2024 Main Conference Waterfall: Scalable Framework for Robust Text Watermarking and Provenance for LLMs

@@ Abstract
Protecting intellectual property (IP) of text such as articles and code is increasingly important, especially as sophisticated attacks become possible, such as paraphrasing by large language models (LLMs) or even unauthorized training of LLMs on copyrighted text to infringe such IP. However, existing text watermarking methods are not robust enough against such attacks nor scalable to millions of users for practical implementation. In this paper, we propose Waterfall, the first training-free framework for robust and scalable text watermarking applicable across multiple text types (e.g., articles, code) and languages supportable by LLMs, for general text and LLM data provenance. Waterfall comprises several key innovations, such as being the first to use LLM as paraphrasers for watermarking along with a novel combination of techniques that are surprisingly effective in achieving robust verifiability and scalability. We empirically demonstrate that Waterfall achieves significantly better scalability, robust verifiability, and computational efficiency compared to SOTA article-text watermarking methods, and showed how it could be directly applied to the watermarking of code. We also demonstrated that Waterfall can be used for LLM data provenance, where the watermarks of LLM training data can be detected in LLM output, allowing for detection of unauthorized use of data for LLM training and potentially enabling model-centric watermarking of open-sourced LLMs which has been a limitation of existing LLM watermarking works. Our code is available at https://github.com/aoi3142/Waterfall.
EMNLP 2024 Main Conference
Waterfall: Scalable Framework for Robust Text Watermarking
and Provenance for LLMs
@@ 1 Introduction
Adding to the challenge is the growing prevalence of generative large language models (LLMs) that may be trained on copyrighted text without permission. To enforce IP rights, content creators need to do data provenance for LLMs, i.e., prove whether their set of work had been used to train 3rd party black-box LLMs. While there have been recent works tackling these problems (Abdelnabi and Fritz, 2021; Zhang et al., 2023), they largely require intervening in the training process of the LLMs. This is unrealistic in practice, as not all LLM service providers may cooperate due to incentive misalignment, and adversaries may use open-source LLMs.
@@
In this paper, we propose Waterfall, the first training-free framework for robust and scalable text watermarking applicable across multiple text types (e.g., articles, code) and languages supportable by LLMs, for general text as well as LLM data provenance. Rather than viewing LLMs as just sources of IP infringement, we introduce the novel perspective of using LLMs’ capabilities to protect existing IP. Though simple, our training-free framework comprises several key innovations such as being the first to use LLM as paraphrasers for watermarking along with a novel combination of techniques that are surprisingly effective in achieving robust verifiability, scalability, and data provenance for LLMs, surpassing state-of-the-art text watermarking methods. In summary, our contributions are as follows:
  1. 1.
    We introduced a formulation of the robust and scalable text watermarking problem setting and lay out a set of desiderata to be satisfied (Section 2).
  2. 2.
    To tackle the challenges arising from these desiderata, we proposed Waterfall comprising novel innovations, including: (a) effective use of LLM paraphrasers to watermark existing text with IP to be protected (Section 3.1); (b) combination of vocab permutation and a new orthogonal watermarking perturbation method in token space, to achieve high scalability and robust verifiability while preserving fidelity (Section 3.3).
  3. 3.
    We conducted comprehensive empirical evaluations, demonstrating that Waterfall achieves significantly better scalability, robust verifiability, and computational efficiency compared to SOTA article-text watermarking methods (Section 4.1), while meeting the desiderata for a variety of applications, including for LLM data provenance of articles (Section 4.3). We also showed how Waterfall could be directly applied to the watermarking of programming code (Section 4.2).
@@ 2 Problem formulation and Desiderata
Watermarking:
Client $i$ uses a watermarking operator $\mathcal{W}(\mu_{i},T_{\text{o}})\rightarrow T_{\text{w}}^{(i)}$ to produce a text $T_{\text{w}}^{(i)}$ that contains watermark $\mu_{i}$, preserves $c$, and can then be used/distributed freely.
Attacks:
There are adversaries who aim to claim the IP in $T_{\text{w}}^{(i)}$ through attacks $\mathcal{A}(T_{\text{w}}^{(i)})\rightarrow T_{\text{sus}}^{(i)}$ that generate their own text $T_{\text{sus}}^{(i)}$ without the watermark $\mu_{i}$ while preserving semantic content $c$. Adversaries do not know $\mu_{i}$ but are able to perform several classes of attacks:
@@
Verification:
Client $i$ can use a verification operator $\mathcal{V}(\mu_{i},T_{\text{sus}})$ to generate a score $q$ indicating the likelihood that $T_{\text{sus}}$ contains $\mu_{i}$. They can then use a setting-specific threshold $\bar{q}$ to classify $T_{\text{sus}}$ as watermarked with $\mu_{i}$ if $q\geq\bar{q}$. The operator $\mathcal{V}$ should be quick and not assume access to $T_{\text{o}}$, as in practice client $i$ may have a large set of $T_{\text{w}}$ and would need to automate the application of $\mathcal{V}$ to scan through a large set of $\{T_{\text{sus}}\}$ to identify any plagiarism (further discussion in Appendix K).
@@
1. Fidelity.
The watermarked text $T_{\text{w}}$ should be semantically similar to $T_{\text{o}}$, i.e., $\mathcal{S}(T_{\text{o}},T_{\text{w}})\geq s$, where $\mathcal{S}:\mathbb{T}\crossproduct\mathbb{T}\rightarrow[0,1]$ is a user-defined fidelity metric depending on the purpose and type of text (e.g., semantic similarity score for articles, or unit tests for code) and $s$ is a setting-specific threshold. We define $\mathbb{T}_{c,s}^{\mathcal{W}}=\{T\in\mathbb{T}:\mathcal{S}(T_{\text{o}},T)\geq s\}$ as the support set of all $T_{\text{w}}$ that a watermarking operator $\mathcal{W}$ can possibly generate for $T_{\text{o}}$ with content $c$ under a $s$-fidelity setting.
2. Verifiability.
The verification operator $\mathcal{V}(\mu_{i},T_{\text{w}}^{(i)})$ should have high efficacy, accounting for Type I and II errors over various thresholds $\bar{q}$. We evaluate this with AUROC computed over a test set.
@@
3. Robust verifiability.
The verification operator on watermarked text after attacks $\mathcal{A}\in\mathbb{A}$, i.e., $\mathcal{V}(\mu_{i},\mathcal{A}(T_{\text{w}}^{(i)}))$, retains high verifiability. This means that the watermark should remain verifiable even after attacks, which constrains framework design. For example, the verification operator should not extract $\mu$ in any subroutine, as an attacker may use it to get $\mu$ and devise an $\mathbb{A}_{3}$ attack to overwrite it (see Section 4.1).
4. Scalability.
The framework should support a large $|\mathbb{M}|$ (set of IDs) while meeting all other desiderata.
@@ 3.3 Increasing scalability with vocab permutation and orthogonal perturbation
Orthogonal perturbation:
Our proposed perturbation operator $\mathcal{F}$ involves two sub-operations acting on $V_{w}$. It first maps each key $k_{p}\in\mathbb{K}_{p}$ to a unique function in a pre-defined family of orthogonal functions, and then adds the chosen perturbation function to the logits $L_{j}$ of the LLM output in $V_{w}$ space:
@@ 3.4 Waterfall Framework
Practical considerations. Waterfall is highly adaptable, i.e., it can be implemented with different LLM as paraphrasers, allowing our framework to achieve better watermarking performance and support more text types as the LLM landscape evolves. Methods like prompt engineering (Wei et al., 2022; Lin et al., 2023) and Reflexion (Shinn et al., 2023; Madaan et al., 2023) may also help to boost performance in some settings, as we demonstrate in our code watermarking experiments (Section G.2). Additional engineering techniques applied to the LLM paraphraser generation process could also further improve the performance of Waterfall. For example, using beam search (Shao et al., 2017) instead of just token-level sampling and choosing the best watermarked text version that balances between fidelity and verifiability results in significantly better watermarking, beyond the results presented in Section 4. We elaborate further on this implementation, along with possible large-scale deployment methods of Waterfall and other practical considerations in Appendix L.
@@ 4 Experiments
Figure 4: Higher watermarking strength $\kappa$ improves verifiability and extraction accuracy. (a) Increasing $\kappa$ trades off fidelity for higher verifiability. (b) Waterfall performs significantly better than benchmarks in text quality, achieving a text quality-verifiabilty Pareto frontier that is much higher benchmarks (plotted as single points since they do not have adjustable settings to balance the quality-verifiability trade-off).
Figure 5: (a) Longer token length $N$ improves verifiability. (b) Combining more pieces of text improves extraction accuracy towards $100\%$. Extraction accuracy is significantly higher than random guess accuracy of $0.003125\%$.
@@ 4.1 Data ownership
Fidelity-verifiability.
We consider the fidelity and verifiability of the schemes before adversarial attacks. Verifiability is computed as the AUROC based on varying their respective classification thresholds, i.e., the verification score threshold $\bar{q}$ for Waterfall, and bit-error rate threshold for M-bit and P-nlw.
Waterfall allows for adjustable watermarking strength to calibrate the fidelity-verifiability trade-off based on the clients’ needs. Figure 4a shows the Pareto frontier of the trade-off. Stronger watermark strength $\kappa$ improves verifiability but also introduces larger distortions to the LLM paraphrasing process, decreasing the fidelity of watermarked text. For our experiments, we mainly used $\kappa=6$, achieving a mean AUROC of 0.992 and STS of 0.887. Even for shorter texts of just 100 tokens ($about$ 65 words), Waterfall achieves high verifiability with an AUROC of 0.98 (Figure 4b). Additional results are in Section E.4. Note that M-bit and P-nlw allows for only a single fidelity-verifiability score, with mean STS scores of 0.998 and 0.942 respectively, and corresponding AUROC scores of 0.987 and 0.882. While the STS scores are high, it is expected as the schemes only make minor edits to $T_{\text{o}}$ which would be more fragile to attacks, as we will see later.
Text fluency and quality.
In fact, both M-bit and P-nlw causes significant degradation in text fluency and quality despite their high STS scores, significantly impacting its practicality in real-life applications. For example, the word replacements by M-bit and P-nlw introduced noticeable linguistic errors that are not captured by the STS score (shown in Section E.5). In contrast, Waterfall consistently produce high quality watermarked text. To systematically analyze this, we quantify text quality via evaluations using ChatGPT, which has been shown to be a good evaluation metric for text fluency (Wang et al., 2023a). Details of the evaluation method, along with results using another text quality metric (GPTScore (Fu et al., 2024)) which further corroborate our findings, are in Section E.6. We can see in Figure 4b that Waterfall significantly outperforms benchmark methods for the fluency score across the different watermark strengths while achieving high verifiability (AUROC), and the results of M-bit and P-nlw lie far within the Pareto frontier of Waterfall.
Robust verifiability.
We consider the various classes of attacks $\mathbb{A}$ mentioned in Section 2. Details of the setup for each attack and additional results are in Appendix F.
@@
Scalability.
As mentioned in Section 3.3, Waterfall has a large maximum scalability of $M=|\mathcal{F}_{1}|\cdot|\mathbb{V}|!\sim 10^{130274}$ based on our implementation using the Fourier perturbation function, and Llama-2 model as LLM paraphraser. In comparison, M-bit and P-nlw, have scalability dependent on the number of possible synonym replacements in any given text, which is limited by text length and varies for different text. On the c4 dataset with a mean article length of 355 words, M-bit and P-nlw can only embed a mean of 9.5 bits ($M\sim 10^{3}$) and 23.2 bits ($M\sim 10^{10}$) respectively.
@@
Metadata extraction.
We also demonstrate how Waterfall could be used to embed metadata while watermarking. We consider metadata $k_{p}\in\{1,2,...,31999\}$, and the task is to extract the embedded $k_{p}$ if the text has been verified as watermarked with $\mu$. We do so by using $k_{p}$ as the frequency of the Fourier perturbation function $\mathcal{F}_{1}$, and perform extraction with the DFT. Figure 5b shows the extraction accuracy of Waterfall for different perturbation magnitudes $\kappa$. By taking multiple samples of $T_{\text{w}}$, multiple articles could be combined together to improve extraction accuracy. For $\kappa=6$, accuracy increases from 48% to 99% with only 5 pieces of text. Details are in Section E.9.
Computational costs.
We note that Waterfall also has lower computational cost compared to benchmarks (Table 2). Waterfall verification can be run in parallel on a CPU, requiring only 0.035s when ran on a single 16-core CPU, which is 75x and 4237x faster than M-bit and P-nlw respectively, both which require inference using deep learning models. This is important in the context of protection of IP, e.g., where data providers may have to scan through large amount of online data for any IP infringement. Further discussion on the deployment costs are in Appendix L.
@@ 4.3 LLM data provenance of articles
Figure 7: More queries improve LLM attribution verifiability. Increasing number of clients $M$ in the training dataset only slightly decreases verifiability. Note that the curve for $M=20$ overlap with that of $M=100$.
@@
Verifiability.
To evaluate verifiability, we queried the fine-tuned model with the first 50 tokens of a randomly chosen abstract, and applied the verifiability operator on the next 100 generated new tokens to test for the associated watermark (details in Section H.2). Our results, presented in Figure 7, shows that Waterfall has high verifiability, reaching AUROC of 1.0 with just 100 queries to the fine-tuned LLM.
Scalability.
To explore the scalability of Waterfall for data provenance, we combined the datasets of different number of clients, $M\in\{1,5,10,20,100\}$, each watermarked with their own unique ID $\mu$, and use the combined dataset for fine-tuning the adversarial model. As expected, Figure 7 shows that dealing with an aggregated dataset mixed with a larger $M$ number of different watermarks would result in a decrease in verifiability. However, our results indicate that this decrease leveled off from $M=20$ to $M=100$ and still allow for an AUROC (verifiability) of 1.0 with around 100 queries even for $M=100$, demonstrating the scalability of Waterfall to a sizable number of clients.